OpenStudy (anonymous):

please divide 35wz^5/75wz^3 please show work if possible thanks

OpenStudy (anonymous):

is it (35wz)^5 or 35w(z)^5

OpenStudy (lgbasallote):

\[\large (\frac{35}{75}) \times (\frac{w}{w}) \times (\frac{z^5}{z^3})\]

OpenStudy (lgbasallote):

just separate them like that

OpenStudy (anonymous):

its just  thanks